OpenStudy (ttp):

For the function 3x – 2y = 12, what is the x-intercept? -4 4 6 -6

OpenStudy (solomonzelman):

lets put this into slope intercept form. y=mx+b \[3x – 2y = 12~~~~->~~~~-2y=-3x+12~~~~->~~~~y=(-2/3)x-6\]what is the y intercept?

OpenStudy (solomonzelman):

actually it is positive 2/3

OpenStudy (the_fizicx99):

Y = mx + b 3x -2y = 12 -3x -3x -2y = -3x + 12 /-2 /-2 /-2 y = 3/2x - 6

OpenStudy (the_fizicx99):

@SolomonZelman It's 3/2x

OpenStudy (solomonzelman):

You have to make y=0 and solve for x to find the x intercept.

OpenStudy (solomonzelman):

3x-2(0)=12 x=4

OpenStudy (the_fizicx99):

Not and not negative, 2 negative's cancel out.

OpenStudy (solomonzelman):

I already corrected my self, see above?

OpenStudy (solomonzelman):

3x-2(0)=12 x=4 so it is going to be 4.

OpenStudy (isaiah.feynman):

The x intercept is 4, set y =0 and solve for x. Simple.

OpenStudy (the_fizicx99):

To find any intercept, multiply the other axis by 0 to cancel out. 3x - 2(0) = 12 3x = 12 /3 /3 x = 4 @SolomonZelman You put 2/3 xD

OpenStudy (solomonzelman):

No, I though the question said y-intercept at first, but then showed the work for x-intercept.

OpenStudy (the_fizicx99):

Oh :)

OpenStudy (solomonzelman):

I see the asker got her help; good luck with the rest, @TTP!

OpenStudy (isaiah.feynman):

My medal please?

OpenStudy (the_fizicx99):

LOL ^

OpenStudy (the_fizicx99):

lulz :>